OpenStudy (anonymous):

soybean meal is 12% protein; cornmeal is 6% protein. How many pounds of each should be mixed together in order to get 240lb mixture that is 8% protein? How many pounds of cornmeal should be in the mixture? How many pounds of the soybean meal should be in the mixture?

OpenStudy (anonymous):

i should make a template for this one. seen it maybe 8 times no lie. what book does it come from?

OpenStudy (anonymous):

its off of a final. Im stumped and cannot figure it out.

OpenStudy (anonymous):

8% of 240 is .08*240= 19.2 is as start, so you know you want to end up with 19.2 pounds of protein all together

OpenStudy (anonymous):

lets say x is the number of pounds of soybean, then \[.12x\] is the number of pounds of protein in it

OpenStudy (anonymous):

and if you have x pounds of soybeans, since the total is 240 you know you have 240 - x pounds of cornmeal.

OpenStudy (anonymous):

this is clear i hope. for example if you have 100 soybeans then you have 240 - 100 = 140 pounds of cornmeal

OpenStudy (anonymous):

it contributes \[.06\times (240-x)\] pounds of protein, that is 6% of the number of pounds of cornmeal. so the total pounds of protein is \[.12x+.06(240-x)\] and you want this to be 19.2 so set \[.12x+.06(240-x)=19.2\] and solve for x

OpenStudy (anonymous):

i have been long winded because i hope if you see a different problem you could translate that as well. now you have to solve the equation. do you need help with that? i will be happy to show you the steps

OpenStudy (anonymous):

do you subtract 19.2 to from 240 to get x?

OpenStudy (anonymous):

no lets do the steps to solve for x

OpenStudy (anonymous):

okay. Thank you.

OpenStudy (anonymous):

i don't like decimals much, so i would multiply everything by 100 to remove them. that is, move each decimal over 2 places

OpenStudy (anonymous):

\[12x+6(240-x)=1920\] is that step clear?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

now you have to use the distributive law to remove the parentheses you get \[12x+1440-6x=1920\]

OpenStudy (anonymous):

i multiplied by 6 in the parentheses. now you have to combine like terms on the right: \[6x+1440=1920\]

OpenStudy (anonymous):

if you don't understand a step, please ask

OpenStudy (anonymous):

the last step was \[12x-6x=6x\]

OpenStudy (anonymous):

i understood that part.

OpenStudy (anonymous):

ok so now we have \[6x+1440=1920\] and now you can finally solve for x. subtract 1440 from both sides, then divide by 6. let me know what you get

OpenStudy (anonymous):

80?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

awesome

OpenStudy (anonymous):

at the beginning i wrote "x = number of pounds of soybeans" so you know that the number of pounds of soybeans is 80 and therefore the number of pounds of cornmeal is 240 - 80 = 160. 80 pounds of soybeans, 160 pounds of cornmeal

OpenStudy (anonymous):

one more thing if you are still there.

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

when you start out one of these problems, don't just start by writing an equation in x. state what x is. write it down. that way at the end you know what you have solved for. and don't write "x = soybeans" x is not soybeans, soybeans are soybeans, x is a number. so write "x = the number of pounds of soybeans" or "x= the number of gallons of 4% solution" or whatever. it will really help you

OpenStudy (anonymous):

i hope all the steps are clear enough that you could try one on your own.

OpenStudy (anonymous):

they have been helpful.